OpenStudy (anonymous):

CORRECT ANSWER GETS MEDAL :What is the value of the x variable in the solution to the following system of equations? 3x + 2y = 6 2x + 3y = −1 −4 4 −3 3

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

anyone?

OpenStudy (wolf1728):

A) 3x + 2y = 6 B) 2x + 3y = −1 Multiplying Equation A by -1.5 A) -4.5x -3y = -9 then adding this to B) 2x + 3y = -1 Sums to -2.5 x = -10 x = -10/-2.5 x = 4